find the complete integral and singular integral of 4(1+x3)=9z4pq.


1
Expert's answer
2022-03-19T02:41:41-0400

4(1+x3)=9z2pq

f(x,y,z,p, q)=4x3- 9z2pq+4=0.............(1)


"\frac{dp}{\frac{dF}{dx}+P\frac{dF}{dz}}=\frac{dq}{\frac{dF}{dy}+q\frac{dF}{dz}}=\frac{dz}{-P\frac{dF}{dp}-q\frac{dF}{dq}} 
=\frac{dx}{-\frac{dF}{dx}} =\frac{dy}{-\frac{dF}{dq}}"


"\frac{dp}{12x^2 -18zp^2q}=\frac{dq}{-18zpq^2}=\frac{dz}{9z^2pq-9z^2pq}
=\frac{dx}{-9z^2q}=\frac{dy}{-9z^2p}"

\

Solving this we get dz=0

Integrating (1), we equate q=C1

"4x^3-9z^2pC1+4=0"


"P=\frac{4x^3+4}{9z^2C1}"


"dz= pdx+qdy"


"dz= \frac{4(x^3+4)}{9z^2C1}dx+C1dy"


"dz-C1dy= \frac{4(x^3+1)}{9z^2C1}dx"


"d(z-C1y)=\frac{4(x^3+1)}{9z^2C1}"

Integrating this we get,

"log(z-C1y)=log9z^2C1+log C2"

C2 is an arbitrary constant

"Z-C1y =\frac{4}{9z^2C1}+C2.............(2)"

(2) is our complete integral

To get the singular integration we add 1 and 2

Differentiate 2 with respect to c1 and c2 respectively we get,


"y=9z^2C2"

"0=9z^2c1, c1=0
C2=\frac{y}{9z^2}"

z=0

Therefore y=c2